# Multi-modal Stance Detection: New Datasets and Model

### v1

@@ Abstract
[…]
@@ 1 Introduction
3) A series of experiments on our datasets show that the proposed method significantly outperforms the baseline models [^1] To facilitate future research, all datasets and code are publicly available at url..
@@ 4 Methodology
Dataset Target
Textual Targeted Prompt
Mtse DT
The stance on Donald Trump is:
Mccq CQ
The stance on the use of Chloroquine and Hydroxychloroquine for the treatment or prevention from the coronavirus or COVID 19 is:
Mwtwt CVS_AET
The stance on merger and acquisition between CVS Health and Aetna is:
Mruc RUS
The stance on Russia is:
Mtwq MOC
The stance on Mainland of China is:
@@ 4.4 Multi-modal Stance Detection
TMPT+CoT ${}^{\star}$ ${}^{\star}$ ${}^{\star}$ ${}^{\star}$ ${}^{\star}$ ${}^{\star}$ ${}^{\star}$ 81.10 ${}^{\star}$ ${}^{\star}$ ${}^{\star}$ ${}^{\star}$
@@ 5.3 Experimental Settings
TMPT+CoT 45.58 56.16 ${}^{\star}$ ${}^{\star}$ 67.28 63.78 ${}^{\star}$ ${}^{\star}$ 54.17 48.90
@@ C.1 Analysis of Textual Prompt Tuning
To analyze the impact of the type of the hand-design textual prompts, we design several types of textual prompts and report the experiments in Table 8. Take the MTSE2020 dataset as an example, it can be seen that the experimental results of all types of prompts are superior to those without prompts (BERT+ViT). This demonstrates the significance of the targeted prompts in this task. Further, we can also see that there are considerable differences in performance between different types of prompts. Specifically, when using both target and stance to design prompts (3⃝, 4⃝, and 5⃝), the model performs significantly better. Therefore, in our method, we choose the type of 5⃝ to devise the textual prompts for multi-modal stance detection.
@@
BERT+ViT
63.28
1⃝
Trump
63.32
2⃝
Donald Trump
63.57
3⃝
stance on Donald Trump
64.67
4⃝
What is the stance on Donald Trump?
64.59

### v2

@@ Abstract
Bin Liang1,2,4∗, Ang Li1,2thanks:  The first two authors contribute equally to this work., Jingqian Zhao1,2, Lin Gui5, Min Yang6, Yue Yu3, Kam-Fai Wong4, and Ruifeng Xu1,2,3thanks:  Corresponding Author 1 Harbin Institute of Technology, Shenzhen, China 2 Guangdong Provincial Key Laboratory of Novel Security Intelligence Technologies 3 Peng Cheng Laboratory, Shenzhen, China 4 The Chinese University of Hong Kong, Hong Kong, China  5 King’s College London, UK 6 SIAT, Chinese Academy of Sciences, Shenzhen, China bin.liang@cuhk.edu.hk,  angli@stu.hit.edu.cn,  xuruifeng@hit.edu.cn
@@ 1 Introduction
3) A series of experiments on our datasets show that the proposed method significantly outperforms the baseline models [^1] To facilitate future research, all datasets and code are publicly available at https://github.com/Leon-Francis/Multi-Modal-Stance-Detection.
@@ 4 Methodology
Dataset Target Textual Targeted Prompt
Mtse DT The stance on Donald Trump is:
Mccq CQ The stance on the use of Chloroquine and Hydroxychloroquine for the treatment or prevention from the coronavirus or COVID 19 is:
Mwtwt CVS_AET The stance on merger and acquisition between CVS Health and Aetna is:
Mruc RUS The stance on Russia is:
Mtwq MOC The stance on Mainland of China is:
@@ 4.4 Multi-modal Stance Detection
TMPT+CoT 74.36 76.57 75.36 75.39 69.78 74.39 73.83 81.10 63.81 71.67 77.91 66.74
@@ 5.3 Experimental Settings
TMPT+CoT 45.58 56.16 66.75 68.69 67.28 63.78 52.24 57.40 54.17 48.90
@@ C.1 Analysis of Textual Prompt Tuning
To analyze the impact of the type of the hand-design textual prompts, we design several types of textual prompts and report the experiments in Table 8. Take the MTSE2020 dataset as an example, it can be seen that the experimental results of all types of prompts are superior to those without prompts (BERT+ViT). This demonstrates the significance of the targeted prompts in this task. Further, we can also see that there are considerable differences in performance between different types of prompts. Specifically, when using both target and stance to design prompts (③, ④, and ⑤), the model performs significantly better. Therefore, in our method, we choose the type of ⑤ to devise the textual prompts for multi-modal stance detection.
@@
BERT+ViT - 63.28
Trump 63.32
Donald Trump 63.57
stance on Donald Trump 64.67
What is the stance on Donald Trump? 64.59